What is 6x + 5 = 13 + 2x
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

WHAT IS 6X + 5 = 13 + 2X
----------
It's an equation.

Answer by ewatrrr(24785)

 
Hi,
Note: use of non-caps, much easier to read for most.
6x + 5 = 13 + 2x
4x = 8
x = 2